OpenStudy (anonymous):

27,000 in scientific Notation?

OpenStudy (austinl):

\[2.7\times 10^4\]

OpenStudy (anonymous):

27 x 102

OpenStudy (anonymous):

not 102 but 10 to the power of 2

OpenStudy (austinl):

No

OpenStudy (anonymous):

@austinL I think so.. You can say im wrong but I think Im right.

OpenStudy (anonymous):

@austinL its called opinions

OpenStudy (austinl):

Multiply your answer back out then? First off, that isn't scientific notation. You can only have ones digits and decimals in scientific notation. Secondly, Multiply your answer back out. \[27\times 10^2\] \[10^2 = 100\] \[27\times 100 = 2,700\]

OpenStudy (anonymous):

@austinL one more thing so what is the scientific notation of 13,500

OpenStudy (austinl):

\[1.35\times 10^4\] Do you see what I am doing to arrive at the answer?

OpenStudy (anonymous):

no what are you doing?

OpenStudy (austinl):

you move the decimal to the left until there is one number in front of it. 13,500 You move it to the left and you have one digit in front of the decimal. \[1.35\times10^x\] x is the number of places from the original position you moved it.

OpenStudy (anonymous):

Okay, could you please help me with this too?

OpenStudy (anonymous):

7 x (1 x 109) what does this equal?

OpenStudy (austinl):

is that \[7\times (1\times 10^9)\]

OpenStudy (anonymous):

yes sorry it is that, youre correct

OpenStudy (austinl):

\[10\times 10\times 10\times 10\times 10\times 10\times10\times10\times10\times = ?\]

OpenStudy (anonymous):

1000000000

OpenStudy (austinl):

Correct, now multiply that by 7.

OpenStudy (anonymous):

7000000000

OpenStudy (austinl):

Yep, and for future reference commas are very nice. 7,000,000,000 It makes things much easier to read.

OpenStudy (anonymous):

austin, i need to go help my mother could you please do this for me! :( First, find the number of pills he took on his last trip by counting the number of letters in your first name and multiplying that number by 1 x 10^4. Then triple that amount for this trip (1 point). Write an email to the doctor using the number of pills he needs on this trip written in scientific notation (3 points) and in short word form for a newspaper article headline (3 points). Be sure to include your name at the end.

OpenStudy (anonymous):

I have 5 letters in my name aswell

OpenStudy (austinl):

Ok, I am not supposed to DO problems for you. The scientific notation is one thing. But a fully fledged problem is another.

OpenStudy (anonymous):

so you can help me with this?

OpenStudy (anonymous):

okay could you walk me through it then? @austinL

OpenStudy (austinl):

\[5\times (1\times10^4) = 5\times(10\times10\times10\times10)\]